OpenStudy (anonymous):

Which one is not a valid value of x? √x^2–10x+25+12 √x=15 √x ? how do you solve this? here are the answers a)17.57 b)14.25 c)1.757 d)1.425 i need help! :(

OpenStudy (anonymous):

Just plug in the options you have for x.

OpenStudy (anonymous):

Then get rid of which one is not valid.

OpenStudy (anonymous):

i got 17.57?

OpenStudy (anonymous):

im not sure i dont want to get it wrong :(((((

OpenStudy (anonymous):

let me see if i can read it

OpenStudy (anonymous):

\[\sqrt{x^2-10x+25}+12\sqrt{x}=15\sqrt{x}\]?

OpenStudy (anonymous):

its weird!

OpenStudy (anonymous):

YEA THATS HOW IT LOOKS LIKE :D

OpenStudy (anonymous):

<(*^*)>

OpenStudy (anonymous):

ok lets see what we can do

OpenStudy (anonymous):

\[\sqrt{x^2-10x+25}+12\sqrt{x}=15\sqrt{x}\] \[\sqrt{(x-5)^2}+12\sqrt{x}=15\sqrt{x}\] \[x-5+12x\sqrt{x}=15\sqrt{x}\] might be a good start

OpenStudy (anonymous):

ok so far?

OpenStudy (anonymous):

then maybe \[x-5=3\sqrt{x}\] after subtracting now we can solve without too much trouble

OpenStudy (anonymous):

square both sides and get \[x^2-10x+25=9x\] then a quadratic equation \[x^2-19+x+25=0\]

OpenStudy (anonymous):

no, make that \[x^2-19x+25=0\]

OpenStudy (anonymous):

i get the following answers using the quadratic formula \[1.42,17.58\] rounded those are the only two

OpenStudy (anonymous):

I chose 17.57 but i got wrong its okay i still love you :'(